# FakeNewsGPT4: Advancing Multimodal Fake News Detection through Knowledge-Augmented LVLMs

### v1

@@ 4.3 Ablation Study
Figure 4: Ablation study of the world knowledge inherent in large vision-language models.
The Effect of the LVLMs Knowledge.
@@
To analyze the impact of world knowledge derived from LVLMs, we compare our method with the common practice of processing fused embeddings in a supervised classification manner Shao et al. (2023a). This variant model (w/o World Knowledge) replaces the Vicuna model in FakeNewsGPT4 with a binary classifier to predict true/fake labels. As shown in Fig. 4, harnessing the inherent knowledge in LVLMs improves an average performance by 9.92 points over the variant model. Furthermore, for target domains from distinct countries exhibiting huge distribution differences, FakeNewsGPT4 yields more significant improvements in 13.26% and 12.15%. This could be attributed to the fact that world knowledge from LVLMs can effectively guide the representation of agnostic instances.
5 Conclusion
@@
Our work presents FakeNewsGPT4, a novel framework that leverages world knowledge from LVLMs and forgery-specific knowledge augmentation to alleviate the domain shift issue in multimodal fake news detection. Two types of critical knowledge for manipulation reasoning are augmented in FakeNewsGPT4. Specifically, the multi-level cross-modal reasoning module is devised to focus on the cross-modal semantic correlations, while the dual-branch fine-grained verification module gathers local spatial information to encode artifact traces. Then, we transform the generated knowledge into refined embeddings for alignment with language space. The candidate answer heuristics and soft prompts are introduced as supplementary inputs to unleash the extensive knowledge of LVLMs. Extensive experiments verify that FakeNewsGPT4 shows superior cross-domain performance compared to the state-of-the-art methods.
Limitations
While the proposed FakeNewsGPT4 framework demonstrates promising results in multimodal fake news detection, there are several limitations to consider. Firstly, we capitalize on the inherent world knowledge derived from large vision-language models (LVLMs), which learn from extensive internet data, to alleviate the domain shift issue. Since the open world is constantly evolving, we argue that the cross-domain performance may be affected by the lack of updates to world knowledge in LVLMs when confronted with future fake news. In the future, we can try to update world knowledge of LVLMs by model editing techniques. Secondly, our current approach only considers two modalities, vision and text, in multimodal fake news detection tasks. The malicious manipulation of other modalities such as audio, also poses a significant threat to people’s daily lives. Future studies can extend our experiments to explore the joint detection of AIGC data in more modal combinations.
Ethics Statement
It is crucial to clarify that the categorization of fake news relies on the source information provided in the VisualNews dataset Liu et al. (2021a). VisualNews dataset consists of image-text pairs collected from multiple online news agencies and is publicly available for research purposes only. All experiments are performed on publicly available datasets which are commonly used in research on multimodal fake news detection. Hence, we do not think of any potential harm or bias that can be derived from our work.
Appendix
@@ Appendix A Contrastive Class Prompts
Fig. 5 presents a detailed list of prompts utilized for implementing the contrastive class prompt which is introduced in Section 3.1.2 of the main text. Following WinCLIP Jeong et al. (2023), we employ the compositional prompt ensemble to generate texts representing natural and unnatural states. Specifically, we curate prompt templates involving a photo of a [c] and a photo of the [c]. A complete prompt can be composed by replacing the token [c] in the template-level prompt with one of the state-level prompts, either from the natural or unnatural states.
Figure 5: Lists of two state-level prompts considered in this paper to construct contrastive class prompts.
Appendix B Verification Loss
Pixel-wise Verification Loss
The manipulated segmentation map $M_{s}$ is utilized to calculate focal loss Lin et al. (2017) and dice loss Milletari et al. (2016) supervised by the grounding mask. In the multimodal fake news detection task, where the majority of regions in fake images remain pristine, employing focal loss can alleviate the issue of class imbalance. The Focal loss is computed as follows:
$\mathcal{L}_{\text{focal}}=-\frac{1}{n}\sum_{i=1}^{n}{(1-p_{i})^{\gamma}{\rm log}(p_{i})},$ (10)
@@ Appendix B Verification Loss
$\mathcal{L}_{\text{dice}}=1-\frac{2\sum_{i=1}^{n}{y_{i}\hat{y}_{i}}}{\sum_{i=1}^{n}{y_{i}^{2}}+\sum_{i=1}^{n}{\hat{y}^{2}_{i}}},$ (11)
@@
Path-level Verification Loss
To regress the predicted bounding box, the aggregated token $u_{{\rm agg}}$ is fed into the BBox Detector $D_{v}$, which comprises two multi-layer perception (MLP) layers. Then we compute the patch-level verification loss by combining normal L1 loss and generalized Intersection over Union (IoU) loss as follows:
$\hat{b}$ $=D_{v}\left(u_{\rm agg}\right),$ (12)
@@ Appendix C Real-world Distribution Divergence
In Fig. 6, we present word clouds to illustrate the distribution divergence in real-world context across regional perspectives and thematic focus. Notably, the BBC predominantly reports British news, encompassing various subjects such as culture and entertainment. Conversely, the Washington Post tends to emphasize American political news. This disparity leads to variations in word usage across distinct domains. For instance, the commonly used words in BBC news include “prime minister", “children", and "family" etc, while in the Washington Post news, prevalent terms are “president", “Donald Trump", and “Obama", etc.
Figure 6: Word clouds of training data sourced from subsets of BBC, The Guardian, USA Today, and The Washington Post, where the size of terms corresponds to the word frequency.
@@ Appendix D Implementation Details
Here we provide detailed implementation details for training on the remaining three subsets: The Guardian, USA TODAY, and The Washington Post. The hyperparameters are the same for all subsets but differentiate the learning rate, batch size, and maximum epoch by dataset scale, detailed in Table 6. Additionally, to ensure a fair comparison, the hyperparameters for the PandaGPT using soft prompt tuning are set to be consistent with the values used in our method. For the state-of-the-art method HAMMER, the hyperparameters are set to default values used in  Shao et al. (2023a).
@@
Table 6: Hyperparameters used in The Guardian, USA TODAY, and The Washington Post subsets.
@@ Appendix E Case Analysis
Fig. 7 and Fig. 8 depict cases from the testing set. The former illustrates comparisons between the vanilla LVLM and FakeNewsGPT4, while the latter showcases cases where either the compared methods or the LVLM using soft prompt tuning predictions are incorrect.
In Fig. 7, Case 1 and Case 2 show cases where the vanilla LVLM model predicts all labels as “Fake” while the FakeNewsGPT4 makes correct predictions. This implies that the vanilla LVLM lacks the necessary knowledge to effectively detect fake news when learning from the general corpus, making it challenging to accurately characterize the concept of forgery and provide precise responses. Consequently, the vanilla LVLM relies heavily on common prompt words, such as “face”, resulting in all labels being predicted as “Fake”.
Figure 7: Cases in the testing set where vanilla LVLM and FakeNewsGPT4 confront with pristine news and fake news. ( ) indicates wrong prediction and ( ) indicates correct prediction.
Figure 8 illustrates comparisons between our method with the state-of-the-art method HAMMER, and PandaGPT using soft prompt tuning (PandaGPT+SPT). In Case 3, the absence of forgery-specific knowledge impedes the ability of PandaGPT to perform manipulation reasoning, especially when confronted with fake images containing subtle artifacts. In Case 4, the powerful model HAMMER struggles to handle agnostic instances due to the lack of prior information, resulting in incorrect judgments regarding unseen fake news. Conversely, Case 5 demonstrates the effectiveness of our FakeNewsGPT4, which integrates inherent world knowledge from LVLMs and incorporates forgery-specific information to make accurate predictions.
Figure 8: Cases in the testing set where at least one in the Baseline and the LVLM using soft prompt tuning made incorrect predictions. ( ) indicates wrong prediction and ( ) indicates correct prediction.

### v2

@@ 4.3. Ablation Study
Figure 3. Ablation study of the world knowledge inherent in large vision-language models.
4.3.3. The Effect of the LVLMs Knowledge.
@@
To analyze the impact of world knowledge derived from LVLMs, we compare our method with the common practice of processing fused embeddings in a supervised classification manner (Shao et al., 2023a). This variant model (w/o World Knowledge) replaces the Vicuna model in FKA-Owl with a binary classifier to predict true/fake labels. As shown in Fig. 3, harnessing the inherent knowledge in LVLMs improves an average performance by 9.92 points over the variant model. Furthermore, for target domains from distinct countries exhibiting huge distribution differences, FKA-Owl yields more significant improvements in 13.26% and 12.15%. This could be attributed to the fact that world knowledge from LVLMs can effectively guide the representation of agnostic instances.
Figure 4. Ablation study of the potential module choice of using pre-trained artifact detector to replace visual-artifact localization module.
4.3.4. Analysis on Potential Module Choice.
We replace the artifact extractor module in our framework for other design choices. This variant model (w/ Pre-trained Artifact Detector) replaces the visual-artifact localization module in FKA-Owl with the off-the-shelf pre-trained detector (Shao et al., 2023a) to obtain the artifact embeddings. The pre-trained detector employs the VIT backbone with supervised training by grounding annotations. As depicted in Fig. 3, our visual-artifact localization module brings a significant improvement over the variant model. This could be attributed to the fact that Our designed module leverages the intrinsic visual encoder in LVLMs to extract artifact traces, thereby alleviating the burden of aligning forgery-specific knowledge with LVLMs.
5. Conclusion
@@
Our work presents FKA-Owl, a novel framework that leverages rich world knowledge from LVLMs and enhances them with forgery-specific knowledge, to tackle the domain shift issue in multimodal fake news detection. Two types of critical forgery-specific knowledge are augmented in FKA-Owl: semantic correlation between text and images and artifact trace in image manipulation. To inject this knowledge into the LVLM, we first propose two lightweight specialized modules to learn their representations respectively. Then, we transform the generated knowledge into refined embeddings for alignment with language space. The candidate answer heuristics and soft prompts are introduced as supplementary inputs to unleash the extensive knowledge of LVLMs. Extensive experiments verify that FKA-Owl shows superior cross-domain performance compared to the state-of-the-art methods.
Acknowledgements.
This research is sponsored by National Natural Science Foundation of China (Grant No. 62306041), Beijing Nova Program (Grant No. Z211100002121106, 20230484488), National Natural Science Foundation of China (No.62176025), National Natural Science Foundation of China (No. U21B2045). This work is partially funded by Youth Innovation Promotion Association CAS (Grant No.2022132), and Beijing Nova Program (20230484276).
Supplementary Material
@@ Appendix A Contrastive Class Prompts
Fig. A5 presents a detailed list of prompts utilized for implementing the contrastive class prompt. Following WinCLIP (Jeong et al., 2023), we employ the compositional prompt ensemble to generate texts representing natural and unnatural states. Specifically, we curate prompt templates involving a photo of a [c] and a photo of the [c]. A complete prompt can be composed by replacing the token [c] in the template-level prompt with one of the state-level prompts, either from the natural or unnatural states.
Figure A5. Lists of two state-level prompts considered in this paper to construct contrastive class prompts.
Appendix B Localization Loss
B.1. Pixel-wise Localization Loss
The manipulated segmentation map $M_{s}$ is utilized to calculate focal loss (Lin et al., 2017) and dice loss (Milletari et al., 2016) supervised by the grounding mask. In the multimodal fake news detection task, where the majority of regions in fake images remain pristine, employing focal loss can alleviate the issue of class imbalance. The Focal loss is computed as follows:
(10) $\mathcal{L}_{\text{focal}}=-\frac{1}{n}\sum_{i=1}^{n}{(1-p_{i})^{\gamma}{\rm log}(p_{i})},$
@@ B.1. Pixel-wise Localization Loss
(11) $\mathcal{L}_{\text{dice}}=1-\frac{2\sum_{i=1}^{n}{y_{i}\hat{y}_{i}}}{\sum_{i=1}^{n}{y_{i}^{2}}+\sum_{i=1}^{n}{\hat{y}^{2}_{i}}},$
@@
B.2. Path-level Localization Loss
To regress the predicted bounding box, the aggregated token $u_{{\rm agg}}$ is fed into the BBox Detector $D_{v}$, which comprises two multi-layer perception (MLP) layers. Then we compute the patch-level Localization loss by combining normal L1 loss and generalized Intersection over Union (IoU) loss as follows:
(12) $\hat{b}$ $=D_{v}\left(u_{\rm agg}\right),$
@@ Appendix C Real-world Distribution Divergence
In Fig. C6, we present word clouds to illustrate the distribution divergence in real-world context across regional perspectives and thematic focus. Notably, the BBC predominantly reports British news, encompassing various subjects such as culture and entertainment. Conversely, the Washington Post tends to emphasize American political news. This disparity leads to variations in word usage across distinct domains. For instance, the commonly used words in BBC news include “prime minister”, “children”, and ”family” etc, while in the Washington Post news, prevalent terms are “president”, “Donald Trump”, and “Obama”, etc.
Figure C6. Word clouds of training data sourced from subsets of BBC, The Guardian, USA Today, and The Washington Post, where the size of terms corresponds to the word frequency.
@@ Appendix D Implementation Details
Here we provide detailed implementation details for training on the remaining three subsets: The Guardian, USA TODAY, and The Washington Post. The hyperparameters are the same for all subsets but differentiate the learning rate, batch size, and maximum epoch by dataset scale, detailed in Table D7. Additionally, to ensure a fair comparison, the hyperparameters for the PandaGPT using soft prompt tuning are set to be consistent with the values used in our method. For the state-of-the-art method HAMMER, the hyperparameters are set to default values used in  (Shao et al., 2023a).
Table D7. Hyperparameters used in The Guardian, USA TODAY, and The Washington Post subsets.
@@
Figure E7. Cases in the testing set where vanilla LVLM and FKA-Owl confront with pristine news and fake news. ( ) indicates wrong prediction and ( ) indicates correct prediction.
@@ Appendix E Case Analysis
Fig. E7 and Fig. E8 depict cases from the testing set. The former illustrates comparisons between the vanilla LVLM and FKA-Owl, while the latter showcases cases where either the compared methods or the LVLM using soft prompt tuning predictions are incorrect.
In Fig. E7, Case 1 and Case 2 show cases where the vanilla LVLM model predicts all labels as “Fake” while the FKA-Owl makes correct predictions. This implies that the vanilla LVLM lacks the necessary knowledge to effectively detect fake news when learning from the general corpus, making it challenging to accurately characterize the concept of forgery and provide precise responses. Consequently, the vanilla LVLM relies heavily on common prompt words, such as “face”, resulting in all labels being predicted as “Fake”.
Figure E8. Cases in the testing set where at least one in the Baseline and the LVLM using soft prompt tuning made incorrect predictions. ( ) indicates wrong prediction and ( ) indicates correct prediction.
Figure E8 illustrates comparisons between our method with the state-of-the-art model HAMMER, and PandaGPT using soft prompt tuning (PandaGPT+SPT). In Case 3, the absence of forgery-specific knowledge impedes the ability of PandaGPT to perform manipulation reasoning, especially when confronted with fake images containing subtle artifacts. In Case 4, the powerful model HAMMER struggles to handle agnostic instances due to the lack of prior information, resulting in incorrect judgments regarding unseen fake news. Conversely, Case 5 demonstrates the effectiveness of our FKA-Owl, which integrates inherent world knowledge from LVLMs and incorporates forgery-specific information to make accurate predictions.